A 1500-Watts electric stove is used to heat 250 g of water (4180 J/(°C kg)). How long will it take for the water to heat up from 25°C to 95°C?
1
Expert's answer
2020-07-20T14:57:46-0400

"Q=P\cdot t=cm\Delta T\to t=\frac{cm\Delta T}{P}=\frac{4180\cdot0.25\cdot(95-25)}{1500}=48.8s"